my question for an assignment is -
demonstrate and explain pythagoras theorem including 3 real life examples from your surroundings.
i measured my phone and it is 9.5cm length, 5cm width and its hypotenuse is 10.5, now i need to show the working out. what is the working out to show this works? thanks :)

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

a^2+b^2=c^2
9.5^2+5^2=c^2
90.25+25=115.25
sq rt 115.2=10.7 cm

Answer by solver91311(24713)



The problem is, you either have an object that doesn't actually have a right angle, and therefore the Pythagorean Theorem doesn't apply (highly unlikely, I think) or you just didn't measure very carefully.

Here's the problem. and . Add them up and you get but . So, redo your measurements and then put the new numbers in to the calculations I just showed you. You'll never get it perfect, but you should, with due care, be able to improve on your measurement error.


John

My calculator said it, I believe it, that settles it